OpenStudy (madmerc):

Solve the equation the square root of the quantity x plus 3 plus 4 equals 5 for the variable. Show each step of your solution process.

OpenStudy (madmerc):

sqrt(x + 3) + 4 = 5 sqrt(x + 3) = 5 - 4 sqrt(x+3) = 1 (sqrt(x + 3))^2 = (1)^2 x + 3 = 1 x = 1 - 3 x = -2

OpenStudy (madmerc):

for flvs algebra 2 exam 2.11